24. On a block of ice, located on a horizontal plane smooth at rest, acts a constant horizontal force of magnitude 40N for 5 sec. (f = 0). As a result of the action of force the block moves 100 m. What is the mass of the ice block?


1
Expert's answer
2021-05-31T08:45:09-0400

The mass of the block can be found from the second Newton's law:


"m = \dfrac{F}{a}"

where "F = 40N" is the force that acts on the block, and "a" is its acceleration.

The distance covered by the block is given by the kinematic equation:


"d = \dfrac{at^2}{2}"

where "d = 100m" is the distance, "t = 5s" is the time. Expressing "a" and substituting it into the first formula, obtain:


"a = \dfrac{2d}{t^2}\\
m = \dfrac{Ft^2}{2d}\\
m = \dfrac{40\cdot 5^2}{2\cdot 100} = 5kg"

Answer. 5kg.